A concrete highway curve of radius 80.0 m is banked at a 13.0 ∘ angle. Part A What is the maximum speed with which a 1200 kg rubber-tired car can take this curve without sliding? (Take the static coefficient of friction of rubber on concrete to be 1.0.) Express your answer with the appropriate units.
Answer #1

max velocity on a banked frictional track is given by

v^2 = rg ( sin x + u cos x) / ( cos x - u sin x)

v^2 = 80* 9.8* ( sin 13 + cos 13) / ( cos 13 - sin 13)

v = 35.42 m/s

=========

Comment in case any doubt, will reply for sure . Goodluck